OpenStudy (anonymous):

Solve the open sentence. –2 ≤ n + 4 ≤ 7 A. n ≥ 6 and n ≤ –4 B. n ≤ 2 and n ≤ 7 C. n ≤ –6 and n ≤ 4 D. n ≥ –6 and n ≤ 3

OpenStudy (texaschic101):

-2 <= n + 4 <= 7 --- subtract 4 -2 - 4 <= n + 4 - 4 <= 7 - 4 -6 <= n <= 3 which is the same as D

OpenStudy (texaschic101):

any questions ?